OpenStudy (anonymous):

Medal/Fan! Help Please A car, a train, a bicycle, and a skateboard are traveling at the same speed. Which object has the greatest kinetic energy? (SC.6.P.11.1) a. car b. train c. bicycle d. skateboard

OpenStudy (anonymous):

@Kainui @TheSmartOne

TheSmartOne (thesmartone):

Do you know what kinetic energy is?

OpenStudy (anonymous):

Its energy that a body posses by virtue of being in motion

TheSmartOne (thesmartone):

Do you know the formula for kinetic energy?

OpenStudy (anonymous):

No

TheSmartOne (thesmartone):

well the formula is \(\Large\rm Kinetic~Energy = \frac{1}{2} \times m \times v^2\) \(\Large\rm Where:\\\ M~ is ~Mass \\\ V ~ is~ Velovity\)

TheSmartOne (thesmartone):

Velocity**********

OpenStudy (anonymous):

Oh so its the skateboard?

TheSmartOne (thesmartone):

Before we jump to conclusions.. :) We know that Velocity is a vector quantity and has both speed and direction. What is something that is similar to velocity but is a scalar quantity? A scalar quantity is something that has only speed. And no direction. I have said the word twice so far ;)

OpenStudy (anonymous):

Is it scalar quantity?

TheSmartOne (thesmartone):

No I want to know a word that is the same thing like Velocity but is a scalar quantity.

TheSmartOne (thesmartone):

The word is in your original question too :)

OpenStudy (anonymous):

Speed?

TheSmartOne (thesmartone):

Exactly!

TheSmartOne (thesmartone):

and what does the question say about speed?

OpenStudy (anonymous):

That there all going at the same speed

TheSmartOne (thesmartone):

Exactly! So if they are all going at the same speed, then the letter v in the formula is a constant! \(\Large\rm Kinetic~Energy = \frac{1}{2} \times m \times v^2\)

OpenStudy (anonymous):

how do we figure out what v^2 is

TheSmartOne (thesmartone):

We don't need to as velocity is the same thing as speed. And that v^2 will be the same thing for all 4 vehicles.

TheSmartOne (thesmartone):

Now the only thing we need to worry about is the m. Which of the 4 objects weighs the most?

OpenStudy (anonymous):

The train

TheSmartOne (thesmartone):

Exactly!

TheSmartOne (thesmartone):

And that will be your answer!

OpenStudy (anonymous):

Oh thank you for your help @TheSmartOne X)

TheSmartOne (thesmartone):

Thank you for participating! @Hisokasbae OpenStudy needs more users like you, who participate in their question so that they can learn something! Keep on asking and answering questions! ~TheSmartOne

TheSmartOne (thesmartone):

I am glad that I was able to help you :) You can tag me next time or @iambatman . He is the actual Physics master :)